Jackson purchases a 10-year 1000 par bond with 8% annual coupons for L to yield 9% annually. Six years later, after the coupon, he sells the bond for G when annual yields are 7%. Find L – G without excel.
Answer #1

Number of Years =10
Par Value =1000
Annual Coupon =8%*1000 =80
YTM =9%
Price for L =PV of Coupons+PV of Par Vaue =80*((1-(1+9%)^-10)/9%+1000/(1+9%)^10 =935.8234

After 6 years number of years left=4
YTM =7%
Price for G =PV of Coupons+PV of Par Vaue =80*((1-(1+7%)^-10)/7%+1000/(1+7%)^10 =1070.24